# No Dimensional Sampling Coresets for Classification

### v1

@@ 1.2 Our Contributions
1.3 Outline
So far, our setting has been restricted to the Euclidean space domain. In Section 2, we generalize the use of Euclidean inner product with reproducing kernels, and thus map the framework to reproducing kernel Hilbert spaces. Subsequently, we introduce the kernelized version of Theorem 5 (Theorem 10), along with a summary of its consequential results in Table 3. This also allows us to reproduce the no dimensional sampling result for kernel density estimates in Theorem 12. Section 3 is devoted to the proof of our main results depending on Radamacher complexity (Theorem 2), and the proof of the kernelized generalization to monotonic coresets in Theorem 5. Then Section 4 studies specific choices for the function $\phi$, such as sigmoid, logistic, svm, and ReLU functions, individually demonstrating the applicability of our result (Theorem 10) for each. Consequently, the proofs of all the results listed in Table 2 will be provided in this section. Section 5 contains the generalized and (in our view) simplified proof of Theorem 3, which gives the classic sensitivity sampling bound using VC dimension. Finally, we close the paper with a section discussing related works, followed by a concluding section.
2 Generalized Framework: Reproducing Kernel Hilbert Space
We can establish the validity of our findings in a broader framework by employing reproducing kernels. In particular, the above-mentioned results are automatically derived when considering $\mathcal{X}=\mathbb{R}^{d}$ with the linear kernel $\boldsymbol{K}(x,w)=\langle x,w\rangle$ as a reproducing kernel Hilbert space $H$. For those less familiar with reproducing kernel Hilbert spaces, maintaining this latter setting throughout the entire paper is helpful.
2.1 Kernelization of the Results
The kernel method is known as a powerful technique in machine learning that enables the handling of non-linear relationships in data by implicitly mapping it to a higher-dimensional space. This flexibility and efficiency make it a widely used approach, especially in scenarios where linear methods may not be sufficient. Reproducing kernel Hilbert spaces (RKHS) are intimately connected to kernel methods. An RKHS is a type of Hilbert space associated with a positive definite kernel function. In the context of machine learning, we can think that the input space is mapped into an RKHS through a feature map. Broadly, the RKHS framework provides a mathematical foundation for understanding how non-linear relationships in data can be effectively captured through implicit mappings into higher-dimensional spaces. The key insight in the context of kernel methods is that the kernel function implicitly represents the inner product in the Hilbert space, for more see (Mercer, 1909; Aronszajn, 1950; Boser et al., 1992; Steinwart and Christmann, 2008; Schölkopf and Smola, 2018). The following definition extends Definition 4 to the context of RKHD cases.
Definition 8
Let $H$ be a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ a probability measure over $\mathcal{X}$, $\phi\colon\mathbb{R}\longrightarrow[0,\infty)$ a non-increasing $L$-Lipschitz function, and $k>0$ a constant. We say $P$ is well-behaved (w.r.t $\phi$ and $\boldsymbol{K}$) if

### v2

@@ 1.2 Our Contributions
2 Proofs of Main Results
2.1 Proof of Theorem 1.1
The proof follows by McDiarmid’s inequality (see lemma C.1), which is a strong concentration inequality that bounds the difference between the sampled mean and the true mean of a function satisfying the bounded differences property: the effect of changing a single observation. Here we outline the proof of Theorem 1.1; full proof in Appendix C.1.
Sketch of proof of Theorem 1.1..
In view of Equation (3), we need to show that, for any iid sample $x_{1},\ldots,x_{m}\in\mathcal{X}$ according to the sensitivity normalized distribution $q=p(x)\frac{s(x)}{S}$, with probability at least $1-2\exp\left(-{\frac{2mt^{2}}{S^{2}}}\right)$,
$\sup_{f\in\mathcal{F}}\Big{|}1-\sum_{i=1}^{m}\frac{T_{f}(x_{i})}{m}\Big{|}\leq 2\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{F}})+t.$
$g(x_{1},\ldots,x_{m})=\sup_{f\in\mathcal{F}}\Big{[}1-\sum_{i=1}^{m}\frac{T_{f}(x_{i})}{m}\Big{]}.$
For each $i\in[m]$ and each $(x_{1},\ldots,x_{m}),(x_{1}^{\prime},\ldots,x_{m}^{\prime})\in\mathcal{X}^{m}$ such that $x^{\prime}_{j}=x_{j}$ for $j\neq i$, observe that
$|g(x_{1},\ldots,x_{m})-g(x_{1}^{\prime},\ldots,x_{m}^{\prime})|\leq\frac{S}{m}.$
Therefore, using McDiarmid’s Inequality, with a probability at least $1-\exp\left(-{\frac{2mt^{2}}{S^{2}}}\right)$, we have
$\sup_{f\in\mathcal{F}}\Big{|}1-\sum_{i=1}^{m}\frac{T_{f}(x_{i})}{m}\Big{|}$ $\leq\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\sup_{f\in\mathcal{F}}\Big{|}1-\sum_{i=1}^{m}\frac{T_{f}(x_{i})}{m}\Big{|}+t$
$\leq 2\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{F}})+t\qed.$
2.2 Proof of Theorem 1.2
This approach to get relative error goes through VC-dimension, which exploits combinatorial properties of range spaces. A $P$-range space is a tuple $(\mathcal{X},P,\mathcal{R})$ where $(\mathcal{X},P)$ is a probability measure space and $\mathcal{R}$ is a subset of $2^{\mathcal{X}}$ whose members are $P$-measurable. Given a $P$-range space, the best relative error possible in general is conditioned on a small additive parameter $\eta>0$ as follows. For an $\varepsilon,\eta\in(0,1)$, the measure $\nu$ on $\mathcal{X}$ is called a relative $(\varepsilon,\eta)$-approximation for $(\mathcal{X},P,\mathcal{R})$ if each $R\in\mathcal{R}$ is $\nu$-measurable and $\left|P(R)-\nu(R)\right|\leq\varepsilon\max(\eta,P(R)).$ Using ideas from Li et al. (2001), then Har-Peled & Sharir (2011) showed that a sufficiently large sample according to $P$ provides a relative $(\varepsilon,\eta)$-approximation for $(\mathcal{X},P,\mathcal{R})$, if the VC-dimension is bounded as $\mathsf{VC}$. Specifically, there is a universal constant $C$ such that, for any $\eta>0$, and $\varepsilon,\delta\in(0,1)$, with probability at least $1-\delta$, iid sample $X=\{x_{1},\ldots,x_{m}\}$ according to $P$ with $m\geq\frac{C}{\eta\varepsilon^{2}}\left(\mathsf{VC}\log\frac{1}{\eta}+\log\frac{1}{\delta}\right)$ satisfies
In order to use this relative error conditioned on $\eta$ to obtain an unconditioned relative error, the key insight is that using an $s$-sensitivity sample for an upper sensitivity function $s$ with total sensitivity $S$, then by setting $\eta=1/S$, we obtain unconditioned relative error. This is formalized in the following in the next lemma – a discrete version of which is implicit in the proof of Theorem 31 in (Feldman et al., 2020).
Lemma 2.1.
With this, the proof of Theorem 1.2 is straightforward.
Proof of Theorem 1.2.
Since $(\mathcal{X},Q,\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{F}},\succ))$ is a $Q$-range, for an $\varepsilon\in(0,1)$ and $\eta=\frac{1}{S}$, Har-Peled & Sharir (2011)’s sampling result implies that there is a universal constant $C$ such that, any iid sample $X=\{x_{1},\ldots,x_{m}\}$ according to $Q$ with $m\geq\frac{4CS}{\varepsilon^{2}}\left(\mathsf{VC}\log S+\log\frac{1}{\delta}\right)$ satisfies
with probability at least $1-\delta$. This implies that the uniform probability measure over $X$ serves as a relative $\left(\frac{\varepsilon}{2},\frac{1}{S}\right)$-approximation for $(\mathcal{X},Q,\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{F}},\succ))$. The proof concludes by using Lemma 2.1 to derive the error co-efficient on the right hand side as $(\frac{\varepsilon}{2}+S\frac{1}{S}\frac{\varepsilon}{2})=\varepsilon$. ∎
We next provide what we believe is a simpler and more direct proof of Lemma 2.1. A key improvement is the use of Fubini’s theorem to transform back and forth between integrating over $\mathcal{X}$ to over the range of the function. We first define some useful shorthand notation: $Q_{\succ}(g,t)=Q(\operatorname{\bf range}(g,\succ,t))=\int_{x\in\mathcal{X}}\mathbbm{1}_{g(x)>t}{\rm d}Q$.
Lemma 2.2.
For any $g=T_{f}\in\mathcal{T}_{\mathcal{F}}$ and $r\geq 0$,
$\int_{x\in\mathcal{X}}\min\{r,g(x)\}{\rm d}Q(x)=\int_{0}^{r}Q_{\succ}(g,t)){\rm d}t.$
Proof.
$\int_{x\in\mathcal{X}}$ $\min\{r,g(x)\}{\rm d}Q(x)$
$=\int_{x\in\mathcal{X}}\left(\int^{r}_{t=0}\mathbbm{1}_{{}_{t<g(x)}}{\rm d}t\right){\rm d}Q(x)$
$=\int^{r}_{t=0}\left(\int_{x\in\mathcal{X}}\mathbbm{1}_{{}_{t<g(x)}}{\rm d}Q(x)\right){\rm d}t\quad\quad\text{(*)}$*)
$=\int_{t=0}^{r}Q_{\succ}(g,t){\rm d}t,$
where ()(*)( ∗ ) is true due to Fubini’s Theorem. ∎
Proof of Lemma 2.1.
Dividing by $P(f)=\int_{\mathcal{X}}f(z){\rm d}P(z)$, we can now restate the goal as proving
$\frac{1}{P(f)}\left|P(f)-\int_{\mathcal{X}}\frac{Sf(x)}{s(x)}{\rm d}\nu(x)\right|\leq\varepsilon(1+S\eta).$
For an arbitrary $g=T_{f}\in\mathcal{T}_{\mathcal{F}}$, using that $\int_{\mathcal{X}}g(x){\rm d}\nu=\frac{1}{P(f)}\int_{\mathcal{X}}\frac{Sf(x)}{s(x)}{\rm d}\nu(x),$ and then $\int_{\mathcal{X}}g(x){\rm d}Q(x)=1$ we can transform the left-hand side
${\frac{1}{P(f)}\left|P(f)-\int_{\mathcal{X}}\frac{Sf(x)}{s(x)}{\rm d}\nu(x)\right|}$
$=\left|1-\int_{x\in\mathcal{X}}g(x){\rm d}\nu(x)\right|$
$=\Big{|}\int_{x\in\mathcal{X}}\min\{S,g(x)\}{\rm d}Q(x)$
$\hskip 56.9055pt-\int_{x\in\mathcal{X}}\min\{S,g(x)\}{\rm d}\nu(x)\Big{|},$ (7)
where the last line follows $\max_{x}g(x)\leq S$. Now applying Lemma 2.2 twice on both $Q$ and $\nu$ we have
@@ 2 Proofs of Main Results
$\leq\int_{t=0}^{S}\left|Q_{\succ}(g,t)-\nu_{\succ}(g,t)\right|{\rm d}t$
$\leq\varepsilon\int_{t=0}^{S}\max\{\eta,Q_{\succ}(g,t)\}{\rm d}t,$ (8)
where the last line follows by $\nu$ being an $(\varepsilon,\eta)$-approximation of $(\mathcal{X},Q,\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{F}},\succ))$. We can then split this integral on $t$ from 00 to $S$ into two parts at a point $r_{g}=\sup\{r\geq 0\colon Q_{\succ}(g,t)\geq\eta\}$. Note that for each $t\in[0,r_{g})$, $Q_{\succ}(g,t)\geq\eta$ and for each $t>r_{g}$, $Q_{\succ}(g,t)<\eta$. Taking these observations into account, we obtain
$\int_{t=0}^{r_{g}}\max\{\eta,Q_{\succ}(g,t)\}{\rm d}t$ $=\int_{t=0}^{r_{g}}Q_{\succ}(g,t){\rm d}t$
(Lemma 2.2) $=\int_{x\in\mathcal{X}}\min\{r_{g},g(x)\}{\rm d}Q(x)$
$\leq\int_{x\in\mathcal{X}}g(x){\rm d}Q(x)=1.$
and
$\int_{t=r_{g}}^{S}\max\{\eta,Q_{\succ}(g,t)\}{\rm d}t=\int^{S}_{t=r_{g}}\eta{\rm d}t\leq S\eta.$
Combining these two parts, we have
(8)$\leq\varepsilon(1+S\eta)$ )
completing the proof. ∎
2.3 Proof of Theorem 1.4
When dealing with Rademacher complexity, we can leverage the helpful lemma known as Talagrand’s Contraction Lemma (refer to Theorem 4.12 in (Ledoux & Talagrand, 1991)). This lemma, in particular, establishes that $\mathcal{R}_{m}(\phi\circ\mathcal{F})\leq L\mathcal{R}_{m}(\mathcal{F})$, where $\phi\colon\mathbb{R}\longrightarrow\mathbb{R}$ is an $L$-Lipschitz function, and $\phi\circ\mathcal{F}=\{\phi\circ f\colon f\in\mathcal{F}\}$. Our requirement extends beyond this lemma to the following generalization (for the proof, see Appendix C.2).
Corollary 2.3.
Let $\boldsymbol{T}\subset\mathbb{R}^{m}$ be a bounded set, and $\varepsilon_{1},\ldots,\varepsilon_{m}>0$. For any $L$-Lipschitz functions $\phi_{i}\colon\mathbb{R}\longrightarrow\mathbb{R}$ with $\phi_{i}(0)=0$ for $i\in[m]$, we have
$\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\boldsymbol{\varepsilon}}$ $\sup_{\boldsymbol{t}\in\boldsymbol{T}}\Big{|}\sum_{i=1}^{m}\varrho_{{}_{i}}\phi_{i}(t_{i})\Big{|}\leq 2L\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\boldsymbol{\varepsilon}}\sup_{\boldsymbol{t}\in\boldsymbol{T}}\Big{|}\sum_{i=1}^{m}\varrho_{{}_{i}}t_{i}\Big{|},$
where $\boldsymbol{\varepsilon}=\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}$.
To establish the proof of Theorem 1.4, we first need to bound the Radamacher complexity.
Lemma 2.4.
For a well-behaved measure $P$ (see Definition 1.3), if $s(\cdot):\mathbb{R}^{d}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}_{\phi,k}$ with total sensitivity $S$, then
$\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{L}_{\phi,k}})\leq C\sqrt{\frac{S}{m}}$
for $C=\left(2LE_{1}+\phi(0)\right)\max\left(4E_{1}k,\frac{1}{E_{2}}\right)+8LkE_{1}+1$.
As the proof is detail-involved, we present a sketch of it here. The complete proof can be found in Appendix C.3.
Sketch of proof..
we start with the two following observations
$\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\Big{\|}\sum_{i=1}^{m}\sigma_{i}\frac{x_{i}}{s(x_{i})}\Big{\|}_{2}\leq\sqrt{\frac{m}{S}}E_{1}$ (9)
$\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\Big{|}\sum_{i=1}^{m}\sigma_{i}\frac{1}{s(x_{i})}\Big{|}\leq\sqrt{\frac{m}{S}}.$
Next, for $\alpha(w)=\frac{S}{\int_{\mathcal{X}}\ell_{\phi,k}(w,x){\rm d}P(x)}$, we deduce that $\alpha(w)\|w\|_{2}^{2}\leq Sk$ and
$\alpha(w)\leq\left\{\begin{array}[]{cc}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)&\|w\|_{2}\leq 1\\ \frac{kS}{2^{2n}}&2^{n}\leq\|w\|_{2}\leq 2^{n+1}.\end{array}\right.$
For $\bar{\phi}(\cdot)=\phi(\cdot)-\phi(0)$, we obtain
$\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{L}_{\phi,k}})\leq$
$\quad\underbrace{\operatorname*{\mathbb{E}}_{\begin{array}[]{c}x_{1:m}\sim q\\ \sigma_{1:m}\sim\{-1,1\}\end{array}}\sup_{w\in\mathcal{X}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|}_{=M_{1}}$
$\quad\quad+\underbrace{\operatorname*{\mathbb{E}}_{\begin{array}[]{c}x_{1:m}\sim q\\ \sigma_{1:m}\sim\{-1,1\}\end{array}}\sup_{w\in\mathcal{X}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\phi(0)}{s(x_{i})}\right|}_{=M_{2}}$
$\quad\quad+\frac{1}{k}\underbrace{\operatorname*{\mathbb{E}}_{\begin{array}[]{c}x_{1:m}\sim q\\ \sigma_{1:m}\sim\{-1,1\}\end{array}}\sup_{w\in\mathcal{X}}\alpha(w)\|w\|_{2}^{2}\left|\sum_{i=1}^{m}\frac{\sigma_{i}}{ms(x_{i})}\right|}_{=N}$
Using the provided upper bound for $\alpha(w)$ and Equation (9), we conclude $M_{1}\leq 2LE_{1}\max\left(2E_{1}k,\frac{1}{E_{2}}\right)\sqrt{\frac{S}{m}}+8LkE_{1}\sqrt{\frac{S}{m}}$ and $M_{2}\leq\sqrt{\frac{S}{m}}\max\left(4E_{1}k,\frac{1}{E_{2}}\right)\phi(0)$. Utilizing the bound $\alpha(w)\|w\|_{2}^{2}\leq Sk$ and Equation (9), we derive $N\leq\sqrt{\frac{S}{m}}k$ which completes the proof. ∎
The proof of Theorem 1.4 now follows immediately from Theorem 1.1 and Lemma 2.4.
2.4 Proof of results in Table 2
We here only outline the proofs for $\phi=\operatorname{{\rm logistic}}$. The proofs for other cases share similarities and are presented in full detail in Appendix D. As these results are yielded from Theorems 1.2 and 1.4, we need to know the upper sensitivity function. We start with a simple observation, providing us with a method to calculate an upper sensitivity function.
Lemma 2.5.
For a probability measure $P$ over $\mathcal{X}$, assume that $\mathcal{W}\subseteq\mathcal{X}$, $\ell:\mathcal{X}\times\mathcal{X}\longrightarrow(0,\infty)$, and $\gamma:[0,\infty)\longrightarrow[0,\infty)$ such that $0<\int_{\mathcal{X}}\gamma(\|x\|_{2}){\rm d}P(x)<\infty$ and
$\gamma(\|x\|_{2})\leq\frac{\ell(x,w)}{\ell(y,w)}\quad\quad\forall x,y\in\mathcal{X},w\in\mathcal{W}.$ (10)
Then $s(y)=\frac{1}{\int_{\mathcal{X}}\gamma(\|x\|_{2}){\rm d}P(x)}$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}_{\mathcal{W}})$, where $\mathcal{L}_{\mathcal{W}}=\{\ell(\cdot,w)\colon w\in\mathcal{W}\}$.
Then we need the following extension of a lemma by Tolochinksy et al. (2022).
Lemma 2.6.
$\frac{\phi(-\alpha z)+\frac{z^{2}}{k}}{\phi(\alpha z)+\frac{z^{2}}{k}}\leq\beta(\alpha)\quad 0\leq\alpha\leq B_{1},0\leq z\leq B_{2}.$ (11)
If we set $M=\phi(-B_{1}B_{2})$, then, for each $x,y,w\in\mathcal{X}$ with $\|x\|,\|y\|\leq B_{1},\|w\|\leq B_{2}$, we have
$\frac{\phi(0)}{M\beta(\|x\|)}\leq\frac{\ell_{\phi,k}(x,w)}{\ell_{\phi,k}(y,w)}.$
If $\phi$ is universally bounded by $M$, then we do not need upper bounds $B_{1},B_{2}$ for $\alpha,z$ and consequentially do not need upper bounds for $\|x\|,\|y\|$, and $\|w\|$, and the same statement holds.
With a detailed proof (see Lemma D.8), we observe that
$\frac{\operatorname{{\rm logistic}}(-\alpha z)+\frac{z^{2}}{k}}{\operatorname{{\rm logistic}}(\alpha z)+\frac{z^{2}}{k}}\leq\left\{\begin{array}[]{cc}\frac{85\alpha^{2}k}{\log(\alpha^{2}k)}&\alpha^{2}k>e\\[5.69046pt] 85&\alpha^{2}k\leq e,\end{array}\right.$
for each $\alpha,z\geq 0.$ Using Lemmas 2.5 and 2.6, this concludes $s(y)=1+\frac{170(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ is an upper sensitivity for $(\mathbb{R}^{d},\ P,\ \mathcal{L}_{\operatorname{{\rm logistic}},k})$ with probability measure $P$ with $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{2}\geq A\right\}\right)=0$ (see Lemma D.9).
Theorem 2.7.
Assume that $P$ is probability measure over $\mathbb{R}^{d}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{2}\geq A\right\}\right)=0$, $S=1+\frac{170(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$, and $C=(2A+1)\max(4Ak,2.5)+8Ak+1$. For $m\geq\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}$ from $P$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathbb{R}^{d},P,\mathbb{R}^{d},\ell_{\operatorname{{\rm logistic}},k})$ with probability at least $1-\delta$.
Proof.
We observed that $s(y)=1+\frac{170(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ serves as an upper sensitivity function for $(\mathbb{R}^{d},\ P,\ \mathcal{L}_{\operatorname{{\rm logistic}},k})$. Since it is a constant function, $s$-sensitivity sampling is equivalent to sampling according to $P$. It is worth noting that $\operatorname{{\rm logistic}}$ is a $1$-Lipschitz, convex, and decreasing function. For $E_{1}=A$, we have
$\operatorname*{\mathbb{E}}\limits_{x\sim p}\operatorname{{\rm logistic}}\left(\frac{\|x\|_{2}}{2E_{1}}\right)\geq\operatorname{{\rm logistic}}\left(\frac{1}{2}\right)\geq\frac{2}{5}=E_{2}.$
This implies that Definition 1.3 is satisfied for $\mathcal{L}_{\operatorname{{\rm logistic}},k}$ and thus Theorem 1.4 concludes the statement for $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta}\right)$. ∎
As the upper sensitivity function for $\mathcal{L}_{\operatorname{{\rm logistic}},k}$ is constant, $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)$: every function in $\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}}$ is a function in $\mathcal{L}_{\operatorname{{\rm logistic}},k}$ scaled by a positive constant. Thus ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}},\succ))={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)).$ For an $f_{w}\in\mathcal{L}_{\operatorname{{\rm logistic}},k}$ and $r\geq 0$,
$\operatorname{\bf range}($ $f_{w},\succ,r)=\left\{x\in\mathbb{R}^{d}\colon f_{w}(x)>r\right\}$
=$=\left\{\begin{array}[]{ll}\mathbb{R}^{d}&t\leq 0\\ \left\{x\in\mathbb{R}^{d}\colon\langle x,w\rangle<\log(e^{t}-1)\right\}&t>0,\end{array}\right.$
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)$ only includes half-spaces and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ))\leq d+1$. Leveraging Theorem 1.2, we obtain the following theorem.
Theorem 2.8.
Let $P$ be a probability measure over $\mathbb{R}^{d}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{2}\geq A\right\}\right)=0$ and $S=1+\frac{170(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$. There is an $m=O\left(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right)\right)$ such that any iid sample $x_{1},\ldots,x_{m}$ from $P$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathbb{R}^{d},\ P,\ \mathbb{R}^{d},\ \ell_{\operatorname{{\rm logistic}},k})$ with probability at least $1-\delta$.
3 Conclusion and Experimental results
The appendix shows how these results can be applied to kernelized versions of these problems, and also recover the best sampling bounds for KDE coresets.
Acknowledgements
Impact Statement
This paper advances the field of Machine Learning, particularly in understanding the convergence rate and potential compression of classic problem formulations. While this has potential for a wide variety of impacts, including societal ones, there are none in particular that we feel demand to be highlighted here as specific consequences.
Appendix A Generalized Framework: Reproducing kernel Hilbert Space
We can establish the validity of our findings in a broader framework by employing reproducing kernels. In particular, the above-mentioned results are automatically derived when considering $\mathcal{X}=\mathbb{R}^{d}$ with the linear kernel $\mathsf{K}(x,w)=\langle x,w\rangle$ as a reproducing kernel Hilbert space $H$. For those less familiar with reproducing kernel Hilbert spaces, maintaining this latter setting throughout the entire paper is helpful.
A.1 Kernelization of the Results
The kernel method is known as a powerful technique in machine learning that enables the handling of non-linear relationships in data by implicitly mapping it to a higher-dimensional space. This flexibility and efficiency make it a widely used approach, especially in scenarios where linear methods may not be sufficient. Reproducing kernel Hilbert spaces (RKHS) are intimately connected to kernel methods. An RKHS is a type of Hilbert space associated with a positive definite kernel function. In the context of machine learning, we can think that the input space is mapped into an RKHS through a feature map. Broadly, the RKHS framework provides a mathematical foundation for understanding how non-linear relationships in data can be effectively captured through implicit mappings into higher-dimensional spaces. The key insight in the context of kernel methods is that the kernel function implicitly represents the inner product in the Hilbert space, for more see (Mercer, 1909; Aronszajn, 1950; Boser et al., 1992; Steinwart & Christmann, 2008; Schölkopf & Smola, 2018). The following definition extends Definition 1.3 to the context of RKHS cases.
Definition A.1.
Let $H$ be a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ a probability measure over $\mathcal{X}$, $\phi\colon\mathbb{R}\longrightarrow[0,\infty)$ a non-increasing $L$-Lipschitz function, and $k>0$ a constant. We say $P$ is well-behaved (w.r.t $\phi$ and $\mathsf{K}$) if
$\operatorname*{\mathbb{E}}\limits_{x\sim P}\mathsf{K}(x,x)\leq E^{2}_{1}\qquad\qquad\text{and}\qquad\qquad\operatorname*{\mathbb{E}}\limits_{x\sim P}\phi\left(\frac{\sqrt{\mathsf{K}(x,x)}}{2E_{1}}\right)\geq E_{2},$ (12)